Given the average bond energies in kJ/mol: H-H = 436; Cl-Cl = 242; H-Cl = 431.


What is the energy change for the decomposition of hydrogen chloride? 2HCl -> H2 + Cl2



Is the decomposition of hydrogen chloride an exothermic or endothermic process? How do you know?



1
Expert's answer
2021-05-27T07:40:42-0400

Solution:

Balanced chemical equation:

2HCl → H2 + Cl2


To find enthalpy of the decomposition of hydrogen chloride, use the following formula:

ΔHrxn = ∑(bonds broken) − ∑(bonds formed)

 

In this reaction, 2 H-Cl bonds must be broken. Also, 1 H-H bond and 1 Cl-Cl bond are formed.

 

Bond energies of these bonds:

  • H-Cl bond: 431 kJ/mol
  • H-H bond: 436 kJ/mol
  • Cl-Cl bond: 242 kJ/mol

 

Bonds broken = 2 × (H-Cl) = 2 × 431 kJ/mol = 862 kJ/mol

Bonds formed = 1 × (H-H) + 1 × (Cl-Cl) = 436 kJ/mol + 242 kJ/mol = 678 kJ/mol


ΔHrxn = Bonds broken - Bonds formed = 862 kJ/mol - 678 kJ/mol = 184 kJ/mol

ΔHrxn = +184 kJ/mol

ΔH > 0 - endothermic process


Answer:

ΔHrxn = +184 kJ/mol

ΔH > 0 - endothermic process